# Structural-Spectral Graph Convolution with Evidential Edge Learning for Hyperspectral Image Clustering

### v3

@@ V Conclusion
This paper presents a superpixel clustering method for HSI, centered on two key innovations: a structural-spectral graph convolutional operator (SSGCO) for enhanced HSI superpixel representation and an evidence guided adaptive edge learning (EGAEL) module for superpixel topological graph refinement. SSGCO infuses dedicated spectral convolution into the information aggregation of GNNs, enabling comprehensive extraction of both spatial and spectral information from superpixels. Complementarily, EGAEL adaptively optimizes the superpixel topological graph by learning edge weights under the guidance of empirical edge weights, effectively mitigating the influence of initial incorrect connections. Finally, we integrate the proposed methods into a contrastive clustering framework, leveraging sample alignment and prototype contrast as clustering-oriented training objectives to achieve clustering. Extensive experiments on several HSI datasets demonstrate the effectiveness and superiority of our method.

### v4

@@ V Conclusion
This paper presents a superpixel clustering method for HSI, centered on two key innovations: a structural-spectral graph convolutional operator (SSGCO) for enhanced HSI superpixel representation and an evidence guided adaptive edge learning (EGAEL) module for superpixel topological graph refinement. SSGCO infuses dedicated spectral convolution into the information aggregation of GNNs, enabling comprehensive extraction of both spatial and spectral information from superpixels. Complementarily, EGAEL adaptively optimizes the superpixel topological graph by learning edge weights under the guidance of empirical edge weights, effectively mitigating the influence of initial incorrect connections. Finally, we integrate the proposed method into a contrastive clustering framework, leveraging sample alignment and prototype contrast as clustering-oriented training objectives to achieve clustering. Extensive experiments on several HSI datasets demonstrate the effectiveness and superiority of our method.
@@
[Uncaptioned image] Jianhan Qi received the B.S. degree in Computer Science and Technology from Wuhan University of Technology, Wuhan, China, in 2022, and the M.S. degree in Software Engineering from Southeast University, Nanjing, China, in 2025. His research interests include unsupervised learning and graph neural networks.
[Uncaptioned image] Yuheng Jia (Member, IEEE) received the B.S. degree in automation and the M.S. degree in control theory and engineering from Zhengzhou University, Zhengzhou, China, in 2012 and 2015, respectively, and the Ph.D. degree in computer science from the City University of Hong Kong, Hong Kong, China, in 2019. He is currently an Associate Professor with the School of Computer Science and Engineering, Southeast University, Nanjing, China. His research interests include machine learning and data representation, such as un/semi/weakly-supervised learning, high-dimensional data modeling and analysis, and low-rank tensor/matrix approximation and factorization.
[Uncaptioned image] Hui Liu received the B.Sc. degree in communication engineering from Central South University, Changsha, China, the M.Eng. degree in computer science from Nanyang Technological University, Singapore, and the Ph.D. degree in computer science from the City University of Hong Kong, Hong Kong, China. From 2014 to 2017, she was a Research Associate with the Maritime Institute, Nanyang Technological University. She is currently an Assistant Professor with the School of Computing Information Sciences, Caritas Institute of Higher Education, Hong Kong, China. Her current research interests include image processing and machine learning.
[Uncaptioned image] Junhui Hou (Senior Member, IEEE) is an Associate Professor with the Department of Computer Science, City University of Hong Kong. He holds a B.Eng. degree in information engineering (Talented Students Program) from the South China University of Technology, Guangzhou, China (2009), an M.Eng. degree in signal and information processing from Northwestern Polytechnical University, Xi’an, China (2012), and a Ph.D. degree from the School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore (2016). His research interests are multi-dimensional visual computing. Dr. Hou received the Early Career Award (3/381) from the Hong Kong Research Grants Council in 2018 and the NSFC Excellent Young Scientists Fund in 2024. He has served or is serving as an Associate Editor for IEEE Transactions on Visualization and Computer Graphics, IEEE Transactions on Image Processing, IEEE Transactions on Multimedia, and IEEE Transactions on Circuits and Systems for Video Technology.